OpenStudy (anonymous):

What are the possible rational zeros of f(x) = x4 + 2x3 – 3x2 – 4x + 18 ?

OpenStudy (anonymous):

± 1, ± 2, ± 3, ± 6, ± 9, ± 18 is it ?

OpenStudy (dan815):

are u doing the method where u sub in one of those numbers and see which one works

OpenStudy (anonymous):

noo i didnt do that

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

Yayyy (:

ganeshie8 (ganeshie8):

good work :)

OpenStudy (jack1):

... you have x^4 and yet you have 6 possible answers...???

ganeshie8 (ganeshie8):

Jack, not 6, there are 12 possible answers. that too oly rational zeroes

OpenStudy (dan815):

12* possible answers actually but ya ^^ what he said

OpenStudy (dan815):

these are the numbers that shes gotta try to sub in to see which one works and then do the whole synthetic division BS i hate it

OpenStudy (dan815):

isnt there a quicker way to find the roots of this equation

ganeshie8 (ganeshie8):

we dont need to find zeroes

OpenStudy (dan815):

just asking in general

ganeshie8 (ganeshie8):

question is not about finding zeroes

ganeshie8 (ganeshie8):

ohk...

OpenStudy (anonymous):

If f(x) = 3x – 1 and g(x) = x2 – 2x – 2, what is f(g(x)) how would i do this ?

ganeshie8 (ganeshie8):

take f(x) replace x wid g(x)

ganeshie8 (ganeshie8):

f(x) = 3x – 1 f(g(x)) = 3(g(x)) - 1

ganeshie8 (ganeshie8):

now replace g(x) wid x2 - 2x - 2

OpenStudy (anonymous):

f(g(x)) = 3(x2 - 2x - 2)) - 1

ganeshie8 (ganeshie8):

f(x) = 3x – 1 f(g(x)) = 3(g(x)) - 1 = 3(x2 - 2x - 2) - 1

ganeshie8 (ganeshie8):

good :) simplify

OpenStudy (jack1):

sorry, get it now, "possible answers", thought that meant solve fo y=0 my bad cheers @ganeshie8 @dan815

ganeshie8 (ganeshie8):

aha jack got u :)

OpenStudy (anonymous):

3x²-6x-6-1

ganeshie8 (ganeshie8):

simplify more

ganeshie8 (ganeshie8):

lol u want to leave -6-1 like that ha

OpenStudy (anonymous):

3x²-6x-7

OpenStudy (anonymous):

lol right now ?

ganeshie8 (ganeshie8):

looks good :)

OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = –x6 – x5– x4 – 4x3 – 12x2 + 12 ?

ganeshie8 (ganeshie8):

Descartes rules of signs, heard of that right ?

OpenStudy (anonymous):

Yess. but im not sure how to do it .

ganeshie8 (ganeshie8):

f(x) = –x6 – x5– x4 – 4x3 – 12x2 + 12 step1 : find number of sign changes in f(x)

OpenStudy (anonymous):

1 ?

ganeshie8 (ganeshie8):

Yes, so number of positive zeroes = 1 ------------------(1)

ganeshie8 (ganeshie8):

step2 : find number of sign changes in f(\(\color{red}{-}\)x)

ganeshie8 (ganeshie8):

f(x) = –x6 – x5– x4 – 4x3 – 12x2 + 12 f(-x) = ?

OpenStudy (anonymous):

umm 6 ? im not sure how to do that part.

ganeshie8 (ganeshie8):

first find f(-x) and then we can simply count

ganeshie8 (ganeshie8):

f(x) = –x6 – x5– x4 – 4x3 – 12x2 + 12 f(-x)= -x6 + x5 -x4 + 4x3 - 12x2 + 12

ganeshie8 (ganeshie8):

how many sign changes in f(-x) ?

OpenStudy (anonymous):

ohh ok soo 5 ?

ganeshie8 (ganeshie8):

Yes, 5 changes so, we can have 5, 3, or 1 negative zeroes

OpenStudy (anonymous):

so its Positive: 1; Negative: 5, 3, or 1; Complex: 4, 2, or 0

ganeshie8 (ganeshie8):

so, number of negative zeroes = 5, 3, or 1 ------------------(2)

ganeshie8 (ganeshie8):

brilliant ! thats it !!!

OpenStudy (anonymous):

(;;. Yayyy !

ganeshie8 (ganeshie8):

very good work as usual :)